OpenStudy (yuki):

I need help in linear algebra. It's about finding the transformation matrix. I will put the equation below.

OpenStudy (yuki):

\[T(\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]) = \left[\begin{matrix}10a-9b & 4a+2b \\ -2c+5d & c+2d\end{matrix}\right]\]